OpenStudy (anonymous):

What is the single logarithm of 5(logbM + logbN) - 2 logbC ?

OpenStudy (anonymous):

\[5(\log _{M} + \log _{N}N) - 2 \log _{b}C \]

OpenStudy (loser66):

they are different, which one is correct one?

OpenStudy (anonymous):

Sorry, its, \[5(log _{b}M+log _{b}N)−2log _{b}C\]

OpenStudy (loser66):

\(log _{b}M+log _{b}N=log_{b}MN\) ok?

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

where does the 5 go?

OpenStudy (loser66):

So \(5(log _{b}M+log _{b}N)= log_{b}(MN)^5 =log_{b}M^5N^5\) ok?

OpenStudy (loser66):

and \(2log_{b}C=log_{b}C^2 \) ok?

OpenStudy (loser66):

\[log_{b}M^5N^5-log_{b}C^2=log_{b}{\frac{M^5N^5}{C^2}}\]

OpenStudy (loser66):